### v1

# Quantum resonances and scattering poles of classical rank one locally symmetric spaces

Abstract.
For negatively curved symmetric spaces it is known from [HHP19] that the poles of the scattering matrices defined via the standard intertwining operators for the spherical principal representations of the isometry group are either given as poles of the intertwining operators or as quantum resonances, i.e. poles of the meromorphically continued resolvents of the Laplace-Beltrami operator. We extend this result to classical locally symmetric spaces of negative curvature with convex-cocompact fundamental group using results of Bunke and Olbrich. The method of proof forces us to exclude the spectral parameters corresponding to singular Poisson transforms.
1. Introduction
Let $G$ be a non-compact connected classical real simple Lie group of real rank $1$. We assume for simplicity that the center of $G$ is finite and fix a maximal compact subgroup $K$ of $G$. The quotient $X:=G/K$ is a non-compact Riemannian symmetric space of rank $1$.
Suppose that $\Gamma\subset G$ is a discrete torsion-free subgroup. We denote the limit set of $\Gamma$ in the geodesic boundary $\partial X$ of $X$ by $\Lambda$ and set $\Omega:=\partial X\setminus\Lambda$. Further, we assume that $\Gamma$ is a convex-cocompact subgroup, by which we mean that the quotient $\Gamma\backslash(X\cup\Omega)$ is a compact manifold with boundary. $\Omega$ is then called the domain of discontinuity of $\Gamma$ in $\partial X$ and $\Gamma\backslash\Omega$ is called the boundary at infinity of $\Gamma\backslash X$. We will always assume that $\Gamma$ is not cocompact, which means that the boundary of $\Gamma\backslash(X\cup\Omega)$ is non-empty.
Let $\Delta_{X}:{C^{\infty}_{c}}(X)\to{C^{\infty}_{c}}(X)$ denote the positive Laplacian of $X$ which is a $G$-invariant differential operator. Furthermore, there is the positive Laplacian $\Delta_{\Gamma\backslash X}:{C^{\infty}_{c}}(\Gamma\backslash X)\to{C^{\infty}_{c}}(\Gamma\backslash X)$ of $\Gamma\backslash X$. Since $X$ and $\Gamma\backslash X$ are complete Riemannian manifolds, $\Delta_{X}$ and $\Delta_{\Gamma\backslash X}$ are essentially self-adjoint operators in ${\rm L}^{2}(X)$ and ${\rm L}^{2}(\Gamma\backslash X)$, respectively; we denote their unique self-adjoint extensions again by $\Delta_{X}$ and $\Delta_{\Gamma\backslash X}$.
The quantum resonances are the poles of the meromorphically continued modified ${\rm L}^{2}$-resolvents
$\begin{split}R_{\zeta}&:=(\Delta_{X}-\left\|\rho\right\|^{2}-\zeta^{2})^{-1}:{\rm L}^{2}(X)\to{\rm L}^{2}(X),\\ R^{\Gamma}_{\zeta}&:=(\Delta_{\Gamma\backslash X}-\left\|\rho\right\|^{2}-\zeta^{2})^{-1}:{\rm L}^{2}(\Gamma\backslash X)\to{\rm L}^{2}(\Gamma\backslash X),\end{split}$ (1.1)
[…]
The scattering poles mentioned in the title are the poles of the meromorphic families $S_{\lambda}$ and $S_{\lambda}^{\Gamma}$ of scattering matrices which are constructed from the standard intertwining operators for the spherical principal series representations of $G$. For the definition of the scattering matrices $S_{\lambda}^{\Gamma}$ on $\Gamma\backslash X$ we build on the constructions from [BO00], which for technical reasons exclude the case $G=F_{4}^{-20}$. This is the only reason why we restrict our attention to the classical rank one spaces.
The relation between quantum resonances and scattering poles for the global rank one symmetric spaces was described explicitly in [HHP19]. The quantum resonances are given as the zeros of the function $c(i\zeta)c(-i\zeta)$ with $\mathrm{Im}\,\zeta>0$, where $c$ is a variant of the Harish-Chandra $c$-function. The zeros of $c$ are known (see e.g. [HHP19, Rem. 5.1]). [HHP19, Thm. 7.1]. The main result of that paper says that the scattering poles contained in $\mathcal{H}_{\mathrm{phys}}$ are precisely the quantum resonances, whereas the scattering poles which are not quantum resonances are the poles of the family of standard intertwining operators.
In [Hil24, Problem 6.5] it was suggested that one can use results of Bunke and Olbrich to extend the relation between quantum resonances and scatttering poles to convex-cocompact locally symmetric spaces of rank one. In fact, in [BO12] Bunke and Olbrich studied the meromorphic continuation of the resolvent kernels as well as the scattering poles for the classical locally symmetric spaces with negative curvature and convex-cocompact fundamental group. Their analysis is based on the meromorphic continuation of a family of operators extending $\Gamma$-invariant distributions on $\Omega$ to $\Gamma$-invariant distributions on $\partial X$. Our point of departure is [BO12, Lemma 11], where they show that the difference between the resolvent kernels on $X$ and the lift of the resolvent kernels on $\Gamma\backslash X$ has boundary values on $\partial X\times\partial X$ which depend meromorphically on the spectral parameters. A careful analysis of the proof of this lemma yields a meromorphic function relating the the boundary values of the lifted resolvent kernels of $\Gamma\backslash X$ and the lifted kernels of the scattering matrices. This allows us to generalize the statement about the scattering poles being quantum resonances except for the exceptional spectral parameters for which the Poisson transform is singular so that no boundary value map is defined. More precisely, we can define a physical half-plane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ depending on $\Gamma$ such that our main result can be formulated as follows (see Theorem 5.4).
Theorem. Outside the set $-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}$, the scattering poles which are not in the physical plane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ are precisely the quantum resonances.
Here $\alpha_{0}$ is the unique reduced root in a chosen positive system, see Section 2.
Acknowledgements
BD has received funding from the Deutsche Forschungsgemeinschaft (German Research Foundation, DFG) through the Priority Program (SPP) 2026 “Geometry at Infinity”. JH acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the grant SFB-TRR 358/1 2023 — 491392403.
2. Setup
[…]
Let $A,N\subset G$ be the analytic subgroups with Lie algebras ${\bf\mathfrak{a}}$ and ${\bf\mathfrak{n}}$, respectively, and let $M:=Z_{K}(A)$ be the centralizer of $A$ in $K$. Then the Iwasawa decomposition of $G$ is given by $G=KAN$ and $P:=MAN$ is a minimal parabolic subgroup of $G$. We denote by $\exp:{\bf\mathfrak{g}}\to G$ the exponential map and by $\kappa:G\to K$, $H:G\to{\bf\mathfrak{a}}$ the smooth maps such that $g\in\kappa(g)\exp(H(g))N$ for all $g\in G$. For $Y\in{\bf\mathfrak{a}}$, $a=\exp(Y)\in A$, $\lambda\in{\bf\mathfrak{a}}_{\mathbb{C}}^{*}:={\bf\mathfrak{a}}^{\ast}\otimes_{\mathbb{R}}\mathbb{C}$, we write $a^{\lambda}:=e^{\lambda(Y)}\in\mathbb{C}$. We fix the Haar measure $dk$ on $K$ of volume $1$.
3. Resolvent kernels
It is known from [HC76] and [BO00] that the continuous spectra of $\Delta_{X}$ as well as $\Delta_{\Gamma\backslash X}$ are given by the interval $[\left\|\rho\right\|^{2},\infty)$, with $\Delta_{\Gamma\backslash X}$ having at most finitely many eigenvalues outside this interval. $\left\|\rho\right\|^{2}$ is actually the minimum of the spectrum of $\Delta_{X}$, but in order to describe the minimum of the spectrum of $\Delta_{\Gamma\backslash X}$ we have to introduce the critical exponent
$\begin{split}\delta_{\Gamma}&:=\inf\Big{\{}s>0\,\Big{|}\,\sum_{\gamma\in\Gamma}e^{-sd_{X}(o,\gamma o)}<+\infty\Big{\}}\\ &\phantom{:}=\inf\Big{\{}s>0\,\Big{|}\,\forall\;x,y\in X:\sum_{\gamma\in\Gamma}e^{-sd_{X}(x,\gamma y)}<+\infty\Big{\}},\end{split}$ (3.1)
where $d_{X}:X\times X\to\mathbb{R}$ is the Riemannian distance and $o:=K\in G/K=X$ is the canonical base point. The critical exponent lies in the interval $[0,2\left\|\rho\right\|]$ (see [Web08, Lemma 3.1]).
By [Cor90, Thm. 4.2] (see also [AZ22, Thm. 1] as well as [HC76] and [BO00]) the minimum of the spectrum of $\Delta_{\Gamma\backslash X}$ is $\left\|\rho\right\|^{2}$ if $0\leq\delta_{\Gamma}\leq\left\|\rho\right\|$, whereas in the case $\left\|\rho\right\|<\delta_{\Gamma}\leq 2\left\|\rho\right\|$ the latter minimum is given by $\left\|\rho\right\|^{2}-(\delta_{\Gamma}-\left\|\rho\right\|)^{2}$. Thus, while the family $\zeta\mapsto R_{\zeta}$ is holomorphic on the physical halfplane $\mathcal{H}_{\mathrm{phys}}$, the family $\zeta\mapsto R_{\zeta}^{\Gamma}$ is holomorphic only on the halfplane
$\mathcal{H}^{\Gamma}_{\mathrm{phys}}:=\begin{cases}\{\zeta\in\mathbb{C}\mid\mathrm{Im}\,\zeta>0\}&\text{for }0\leq\delta_{\Gamma}\leq\|\rho\|\\ \{\zeta\in\mathbb{C}\mid\mathrm{Im}\,\zeta>\delta_{\Gamma}-\|\rho\|\}&\text{for }\|\rho\|<\delta_{\Gamma}\leq 2\|\rho\|.\end{cases}$
Denote by $r_{\zeta}\in\mathcal{D}^{\prime}(X\times X)$ and $r^{\Gamma}_{\zeta}\in\mathcal{D}^{\prime}((\Gamma\backslash X)\times(\Gamma\backslash X))$ the Schwartz kernels of $R_{\zeta}$ and $R^{\Gamma}_{\zeta}$, respectively, and by $\tilde{r}^{\Gamma}_{\zeta}\in\mathcal{D}^{\prime}(X\times X)^{\Gamma\times\Gamma}$ the distribution corresponding to $r^{\Gamma}_{\zeta}$ when identifying $\mathcal{D}^{\prime}((\Gamma\backslash X)\times(\Gamma\backslash X))=\mathcal{D}^{\prime}(X\times X)^{\Gamma\times\Gamma}$. Note here that the identification uses the pullback via the canonical projection $\mathrm{pr}_{\Gamma}:X\to\Gamma\backslash X$, which is a submersion (cf. [Hör90, Thm. 6.12]).
When $\zeta$ lies sufficiently deep inside the physical half-plane, $\tilde{r}_{\zeta}^{\Gamma}$ is nothing but the “$\Gamma$-summation” (or “$\Gamma$-average”) of $r_{\zeta}$. To make this precise we formulate the following lemma.
Lemma 3.1 (C.f. [BO12, Lemma 9]).
For all $\zeta\in\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ satisfying $\mathrm{Im}\,\zeta>\sqrt{(\mathrm{Re}\,\zeta)^{2}+c_{\Gamma}}$, one has
$\tilde{r}_{\zeta}^{\Gamma}=\sum_{\gamma\in\Gamma}(e,\gamma)^{\ast}r_{\zeta},$ (3.2)
where
$c_{\Gamma}:=\delta_{\Gamma}^{2}+\max(0,\delta_{\Gamma}-\left\|\rho\right\|)^{2}\in[\delta_{\Gamma}^{2},2\delta_{\Gamma}^{2}].$ (3.3)
and $(\gamma_{1},\gamma_{2})^{\ast}$ denotes the pullback action of an element $(\gamma_{1},\gamma_{2})\in\Gamma\times\Gamma$. In particular, the right-hand side of (3.2) is a well-defined distribution on $X\times X$.
Proof.
We argue as in [BO12, proof of Lemma 9], adding (and repeating) some details. Let $\mathrm{diag}_{X}\subset X\times X$ be the diagonal and define
$D_{\Gamma}:=(\{e\}\times\Gamma)\cdot\mathrm{diag}_{X}=(\Gamma\times\{e\})\cdot\mathrm{diag}_{X}\subset X\times X.$
The proper discontinuity of the $\Gamma$-action on $X$ implies that $D_{\Gamma}$ is a closed subset of $X\times X$ (c.f. [Kap23, Thm. 11 (6)]). As pointed out in [BO12, p. 137], the distribution $r_{\zeta}$ is smooth outside the diagonal, which implies that on $X\times X\setminus D_{\Gamma}$ all the distributions $(e,\gamma)^{\ast}r_{\zeta}$, $\gamma\in\Gamma$, are smooth. Now, let $f\in{C^{\infty}_{c}}(X\times X)$ and let $\Gamma=\{\gamma_{1},\gamma_{2},\ldots\}$ be an enumeration of the group $\Gamma$. Then we have for each $N\in\mathbb{N}$
$\sum_{n=1}^{N}\big{|}((e,\gamma_{n})^{\ast}r_{\zeta})(f)\big{|}=\sum_{\begin{subarray}{c}n\leq N:\\ (\operatorname{supp}f)\cap(e,\gamma_{n}^{-1})\cdot(\operatorname{supp}f)=\emptyset\end{subarray}}\big{|}((e,\gamma_{n})^{\ast}r_{\zeta})(f)\big{|}$
$\qquad\qquad\qquad\qquad\quad+\sum_{\begin{subarray}{c}n\leq N:\\ (\operatorname{supp}f)\cap(e,\gamma_{n}^{-1})\cdot(\operatorname{supp}f)\neq\emptyset\end{subarray}}\big{|}((e,\gamma_{n})^{\ast}r_{\zeta})(f)\big{|}$
$\leq\sum_{\begin{subarray}{c}n\leq N:\\ (\operatorname{supp}f)\cap(e,\gamma_{n}^{-1})\cdot(\operatorname{supp}f)=\emptyset\end{subarray}}\big{|}((e,\gamma_{n})^{\ast}r_{\zeta})(f)\big{|}\;+\sum_{\begin{subarray}{c}\gamma\in\Gamma:\\ (\operatorname{supp}f)\cap(e,\gamma)\cdot(\operatorname{supp}f)\neq\emptyset\end{subarray}}\big{|}((e,\gamma)^{\ast}r_{\zeta})(f)\big{|}.$
Here the last term is a finite sum because the ($(\{e\}\times\Gamma)$-action on $X\times X$ is properly discontinuous. On the other hand, for each $\gamma\in\Gamma$ such that $(\operatorname{supp}f)\cap(e,\gamma^{-1})\cdot(\operatorname{supp}f)=\emptyset$, we can write
$\big{|}((e,\gamma)^{\ast}r_{\zeta})(f)\big{|}$ $=\big{|}((e,\gamma)^{\ast}r_{\zeta})(f|_{X\times X\setminus D_{\Gamma}})\big{|}$
$=\Big{|}\int_{X\times X\setminus D_{\Gamma}}((e,\gamma)^{\ast}r_{\zeta})(x,y)\overline{f(x,y)}\,\mathrm{d}x\mathrm{d}y\Big{|}$
and then apply [BO12, Lemma 8], as well as the expression for the minimum of the spectrum of $\Delta_{\Gamma\backslash X}$ stated above, to estimate for every compact set $W\subset\mathcal{H}^{\Gamma}_{\mathrm{phys}}$:
$\big{|}((e,\gamma)^{\ast}r_{\zeta})(f)\big{|}$ $\leq C\left\|f\right\|_{\infty}e^{-a\sqrt{-\max(0,\delta_{\Gamma}-\left\|\rho\right\|)^{2}-\mathrm{Re}\,(\zeta^{2})}}\quad\forall\;\zeta\in W$
with a constant $C>0$ depending on $W$ and $\operatorname{supp}f$, and the exponent
$a:=\inf_{(x,y)\in\operatorname{supp}f}d_{X}(x,\gamma y)>0.$
Now, in view of the definition of the critical exponent $\delta_{\Gamma}$, the condition for convergence as $N\to\infty$ becomes
$\sqrt{-\max(0,\delta_{\Gamma}-\left\|\rho\right\|)^{2}-\mathrm{Re}\,(\zeta^{2})}>\delta_{\Gamma},$
which is equivalent to
$(\mathrm{Im}\,\zeta)^{2}>(\mathrm{Re}\,\zeta)^{2}+\underbrace{\delta_{\Gamma}^{2}+\max(0,\delta_{\Gamma}-\left\|\rho\right\|)^{2}}_{=c_{\Gamma}}.$
The convergence estimate depends on $f$ only via $\operatorname{supp}f$ and $\left\|f\right\|_{\infty}$, so that we obtain distributional convergence of the right-hand side of (3.2) for all $\zeta\in\mathcal{H}_{\mathrm{phys}}$ with $\mathrm{Im}\,\zeta>\sqrt{(\mathrm{Re}\,\zeta)^{2}+c_{\Gamma}}$.
It remains to show that the right-hand side of (3.2) equals $\tilde{r}_{\zeta}^{\Gamma}$, which can be done exactly as in the end of the proof of [BO12, Lemma 9]. ∎
Remark 3.2.
In (3.2) one could equivalently take $\sum_{\gamma\in\Gamma}(\gamma,e)^{\ast}r_{\zeta}$: since $\Delta_{X}$ is $G$-equivariant, so is each of its resolvents $R_{\zeta}$ for $\zeta\in\mathcal{H}_{\mathrm{phys}}$. Thus $r_{\zeta}$ satisfies $(g,g^{-1})\cdot r_{\zeta}=r_{\zeta}$ for all $g\in G$.
By [BO12, Lemma 6], $r_{\zeta}$ and $r_{\zeta}^{\Gamma}$ (equivalently $\tilde{r}_{\zeta}^{\Gamma}$) extend to $\mathbb{C}$ as meromorphic families of distributions. The meromorphic families of operators ${C^{\infty}_{c}}(X)\to\mathcal{D}^{\prime}(X)$ defined by the distribution kernels $r_{\zeta}$ and $\tilde{r}_{\zeta}^{\Gamma}$ will be denoted by $R_{\zeta}$ and $\tilde{R}_{\zeta}^{\Gamma}$, respectively (where $\tilde{R}_{\zeta}^{\Gamma}$ maps actually into $\mathcal{D}^{\prime}(X)^{\Gamma}$ thanks to the $\Gamma\times\Gamma$-invariance of $\tilde{r}_{\zeta}^{\Gamma}$), while we denote the meromorphic family of operators ${C^{\infty}_{c}}(\Gamma\backslash X)\to\mathcal{D}^{\prime}(\Gamma\backslash X)$ with distribution kernels $r_{\zeta}^{\Gamma}$ by $R_{\zeta}^{\Gamma}$. This is compatible with the previous notation in the sense that $R_{\zeta}$ and $R_{\zeta}^{\Gamma}$ are meromorphic extensions of the operator families from (1.1) with domains shrinked to compactly supported smooth functions and codomains enlarged to distributions.
4. Scattering matrices
In this section we keep the setup from Section 2 and follow [BO00] in the definition of $\lambda\mapsto S_{\lambda}^{\Gamma}$ as a meromorphic family of the scattering matrices for $\Gamma\backslash X$. It is our goal to relate the poles of this family, which we call scattering poles, to the resonances of $\Gamma\backslash X$, i.e. the poles of the meromorphically continued family $\zeta\mapsto R^{\Gamma}_{\zeta}$ of resolvents.
4.1. Spherical principal series representations
In order to introduce the scattering matrices we have to review some facts about spherical principal series representations. We warn the reader that there is no fixed standard notation in the literature. As we will have to use spectral information from different sources, we will have to keep track of the different conventions (see Remark 4.2 below).
For $\lambda\in{\bf\mathfrak{a}}_{\mathbb{C}}^{*}$ we define the one-dimensional group representation
$\sigma_{\lambda}:P=MAN\to\mathbb{C}\setminus\{0\},\quad man\mapsto a^{\rho+\lambda},$ (4.1)
and the associated homogeneous line bundle
$V(\sigma_{\lambda}):=G\times_{P}\mathbb{C}\to G/P,\quad[g,z]\mapsto gP,$
where $[g,z]$ is the $P$-orbit of $(g,z)\in G\times\mathbb{C}$ with respect to the right $P$-action given by $(g,z)\cdot p=(gp,\sigma_{\lambda}(p^{-1})z)$. Note that $G$ acts on the total space of $V(\sigma_{\lambda})$ from the left. The space $C^{\infty}(G/P;V(\sigma_{\lambda}))$ of all smooth sections $s:G/P\to V(\sigma_{\lambda})$ can be identified with
$H^{\lambda,\infty}$ :=$:=\{f\in{C^{\infty}}(G)\mid\forall g\in G,p\in P:f(gp)=\sigma_{\lambda}(p)^{-1}f(g)\}$ (4.2)
so that $f\in H^{\lambda,\infty}$ corresponds to the section $s$ of $V(\sigma_{\lambda})$ with $s(gP)=[g,f(g)]$.
The left-regular representation of $G$ on $H^{\lambda,\infty}$ will be denoted by $\pi^{\lambda,\infty}$:
$\big{(}\pi^{\lambda,\infty}(g)f\big{)}(g^{\prime}):=f(g^{-1}g^{\prime}),\qquad f\in H^{\lambda,\infty}.$ (4.3)
It is called (spherical, smooth) principal series representation with parameter $\lambda$.
Due to the Iwasawa decomposition the restriction map
$\begin{split}H^{\lambda,\infty}&\to{C^{\infty}}(K)^{M}:=\{f\in{C^{\infty}}(K)\mid\forall k\in K,m\in M:f(km)=f(k)\}\\ f&\mapsto f|_{K}\end{split}$ (4.4)
is an isomorphism of vector spaces. This allows us to consider $H^{\lambda,\infty}$ and $C^{\infty}(G/P;V(\sigma_{\lambda}))$ as $\lambda$-independent vector spaces. It also shows that $V(\sigma_{\lambda})$ is in fact a trivial line bundle but we will soon pass to its $\Gamma$-quotient, which need no longer be trivial.
We equip the vector space $C^{\infty}(G/P;V(\sigma_{\lambda}))\cong H^{\lambda,\infty}$ with the Fréchet topology such that (4.4) is an isomorphism of topological vector spaces, where ${C^{\infty}}(K)^{M}$ is equipped with the subspace topology inherited from the standard Fréchet topology on ${C^{\infty}}(K)$. This way, we can consider $H^{\lambda,\infty}$ and $C^{\infty}(G/P;V(\sigma_{\lambda}))$ as $\lambda$-independent Fréchet spaces.
By restricting the line bundle $V(\sigma_{\lambda})$ to $\Omega\subset G/P=\partial X$ and passing to the $\Gamma$-quotient, we obtain the locally homogeneous vector bundle
$V_{\Gamma\backslash\Omega}(\sigma_{\lambda}):=\Gamma\backslash V(\sigma_{\lambda})|_{\Omega}$
over the smooth manifold $\Gamma\backslash\Omega$. The vector space $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ corresponds canonically to the space of $\Gamma$-invariant smooth sections of $V(\sigma_{\lambda})|_{\Omega}$, which in turn corresponds canonically to the closed subspace of left-$\Gamma$-invariant functions in $H^{\lambda,\infty}$. Thus, the space $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ inherits a Fréchet topology from $H^{\lambda,\infty}$. In particular, we can consider $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ as a $\lambda$-independent Fréchet space, too.
Definition 4.1.
We call the topological vector space
$C^{-\infty}(G/P;V(\sigma_{\lambda})):=C^{\infty}(G/P;V(\sigma_{-\lambda}))^{\prime},$
where “prime” denotes the topological dual, the space of distributional sections of $V(\sigma_{\lambda})$.
We equip $C^{-\infty}(G/P;V(\sigma_{\lambda}))$ with the dual $G$-representation of $\pi^{-\lambda,\infty}$, which we denote by $\pi^{\lambda,-\infty}$, and write the subspace of $\Gamma$-invariant elements as ${}^{\Gamma}C^{-\infty}(G/P;V(\sigma_{\lambda}))$. The map
$\begin{split}C^{\infty}(G/P;V(\sigma_{\lambda}))\cong H^{\lambda,\infty}&\to C^{-\infty}(G/P;V(\sigma_{\lambda})),\\ f&\mapsto\left\langle f|_{K},\cdot\right\rangle_{{\rm L}^{2}(K)}\end{split}$ (4.5)
is well-defined, injective (since the restriction (4.4) is injective), and continuous. Moreover, the injection (4.5) is $G$-equivariant; for this, the sign change of $\lambda$ in Definition 4.1 is essential. By virtue of this equivariant injection we can view the principal series representation $(\pi^{\lambda,\infty},C^{\infty}(G/P;V(\sigma_{\lambda})))$ as a subrepresentation of $(\pi^{\lambda,-\infty},C^{-\infty}(G/P;V(\sigma_{\lambda})))$. We call the latter distributional principal series representation with parameter $\lambda$.
4.2. Comparing conventions
We will have to keep track of the conventions for meromorphic continuations with respect to complex parameters used in the various sources we rely on. To simplify this task we include a couple of remarks containing the pertinent comparisons.
Remark 4.2.
Our definitions (4.1), (4.2) differ from the one in [BO00, Section 3], [Olb94, § 2.2] and [HHP19, § 2] by the sign of $\lambda$. We chose this sign convention to be consistent with [Kna86, VII, §1]. This is also consistent with the sign convention in [Vog08, Chap. 3]. In all these references the $G$-action is the left regular action. On the other hand, [Wal92, § 10.1] and [KS71] use the right regular version of the principal series representations. Note that the map $f\mapsto\tilde{f}$ with $\tilde{f}(g):=f(g^{-1})$ intertwines left and right regular actions. Up to this intertwiner the sign conventions in [Wal92, § 10.1] and [KS71] agree with the sign conventions in [Kna86, VII, §1]. Finally, we point out that our topologies on $C^{\infty}(G/P;V(\sigma_{\lambda}))$ and $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ are equivalent to those of [BO00, p. 89].
Remark 4.3 (Compatibility with [BO12, Section 6] and [HHP19]).
[…]
~±${\widetilde{\mathbb{C}}}^{\pm}\owns z\mapsto\sqrt{\pm(\left\|\rho\right\|^{2}-z)}\in\mathbb{C}$ (4.6)
are holomorphic, respectively. As Riemann surfaces, both ${\widetilde{\mathbb{C}}}^{+}$ and ${\widetilde{\mathbb{C}}}^{-}$ are isomorphic to the Riemann surface $\widetilde{\mathbb{C}}$ of the complex square root. This allows us to compare points in ${\widetilde{\mathbb{C}}}^{+}$ and ${\widetilde{\mathbb{C}}}^{-}$ by identifying these spaces with $\widetilde{\mathbb{C}}$. To achieve compatibility with [BO12, Section 6] and [HHP19] one finds that for both ${\widetilde{\mathbb{C}}}^{+}$ and $\widetilde{\mathbb{C}}^{-}$ one needs to choose $[\left\|\rho\right\|^{2},\infty)$ as the branch cut. For ${\widetilde{\mathbb{C}}}^{+}$ this corresponds to the “principal square root” branch cut of $\widetilde{\mathbb{C}}$ along $(-\infty,0]$, while for $\widetilde{\mathbb{C}}^{-}$ it corresponds to the “opposite” branch cut of $\widetilde{\mathbb{C}}$ along $[0,\infty)$. We define the physical sheet ~phys~subscriptsuperscript~$\widetilde{\mathbb{C}}^{-}_{\mathrm{phys}}\subset\widetilde{\mathbb{C}}^{-}$ over $\mathbb{C}\setminus[\left\|\rho\right\|^{2},\infty)$ as the sheet which is mapped bi-holomorphically onto $\mathcal{H}_{\mathrm{phys}}$ by the function with the -- sign in (4.6)
~$\widetilde{\mathbb{C}}^{-}_{\mathrm{phys}}\owns z\mapsto\sqrt{z-\left\|\rho\right\|^{2}}=:\zeta=:\zeta(z)\in\mathcal{H}_{\mathrm{phys}},$
providing us with the “physical parameter” $\zeta\in\mathcal{H}_{\mathrm{phys}}$.
Bunke and Olbrich, on the other hand, work with the parameter
$\lambda:=\lambda(z):=\sqrt{\left\|\rho\right\|^{2}-z}$
obtained by considering the $+$ sign in (4.6). For the conversion from $\lambda$ to $\zeta$ and vice-versa, there are the two possibilities $\lambda=\pm i\zeta$, which amount to choosing $\{\mp\mathrm{Re}\,\lambda>0\}\subset\mathbb{C}$ as the physical $\lambda$-half-plane corresponding to the physical $\zeta$-half-plane $\mathcal{H}_{\mathrm{phys}}$. In [BO12, p. 135] Bunke and Olbrich work with the right half-plane but their argument works just as well with the left half-plane. Indeed, there is no preferred choice of the physical sheet in [BO12, Section 6], so we can use their results regardless of which sign we choose in the relation between the parameters $\lambda$ and $\zeta$. The following choice will turn out to be convenient:
$\begin{split}\lambda&:=\lambda(\zeta):=-i\zeta\qquad\forall\;\zeta\in\mathcal{H}_{\mathrm{phys}},\\ \zeta&=\zeta(\lambda)=i\lambda,\quad\;\forall\lambda\in\mathbb{C},\;\mathrm{Re}\,\lambda>0.\end{split}$ (4.7)
4.3. Scattering matrices
We build on the descriptions of scattering matrices given in [BO00] and [HHP19].
Then, according to [BO00, pp. 79–80], using their isometric identification $\mathbb{C}\owns\lambda\leftrightarrow\lambda\frac{\rho}{\left\|\rho\right\|}\in{\bf\mathfrak{a}}_{\mathbb{C}}^{\ast}$, we have the three meromorphic families on $\mathbb{C}$ of maps
$\mathrm{ext}_{\lambda}$ ::: $C^{-\infty}(\Gamma\backslash\Omega,V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))\to{}^{\Gamma}C^{-\infty}(\partial X,V(\sigma_{\lambda}))$ (4.8)
$\mathrm{res}_{\lambda}$ ::: ${}^{\Gamma}C^{-\infty}(\partial X,V(\sigma_{\lambda}))\mapsto C^{-\infty}(\Gamma\backslash\Omega,V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ (4.9)
$\hat{J}_{\lambda}$ ::: $C^{-\infty}(\partial X,V(\sigma_{\lambda}))\to C^{-\infty}(\partial X,V(\sigma_{\lambda})),$ (4.10)
where the latter is the “unnormalized” Knapp-Stein intertwiner described in detail in [KS71, Thm. 3] (in which the identification $\mathbb{C}\owns\lambda\leftrightarrow\lambda\rho\in{\bf\mathfrak{a}}_{\mathbb{C}}^{\ast}$ is used, so that our $\lambda$ corresponds to $\frac{\lambda}{\left\|\rho\right\|}$ in [KS71, Thm. 3]). Recall from Remark 4.2 that the $P$-representation $\sigma_{\lambda}$ defined in (4.1) is the same as in [KS71] but corresponds to $\sigma_{-\lambda}$ in [BO00]. The Harish-Chandra $c$-function is the meromorphic function on ${\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}$ satisfying
$c(\lambda\rho):=\int_{\bar{N}}e^{-(\lambda+1)\rho(H(\bar{n}))}\mathrm{d}\bar{n},\qquad\forall\;\lambda\in\mathbb{C},\;\mathrm{Re}\,\lambda>0.$
In view of the chosen identification of ${\bf\mathfrak{a}}_{\mathbb{C}}^{\ast}$ with $\mathbb{C}$ and the sign difference in the conventions for the principal series parameter used here and in [BO00], respectively, it is convenient to consider the meromorphic function $c_{\mathbb{C}}$ on $\mathbb{C}$ defined by
$c_{\mathbb{C}}(\lambda):=c\Big{(}\frac{-\lambda}{\left\|\rho\right\|}\rho\Big{)}.$
Then we define the normalized Knapp-Stein intertwiner
$J_{\lambda}:=\frac{1}{c_{\mathbb{C}}(\lambda)}\hat{J}_{\lambda}.$
Note that our $J_{\lambda}$ corresponds to $J_{\mathbf{1},-\lambda}$ in [BO00, p. 101].
Recall the critical exponent $\delta_{\Gamma}$ from (3.1) and note that if $\delta_{\Gamma}^{\mathrm{BO}}$ denotes the critical exponent from [BO00, Def. 2.2], then $\delta_{\Gamma}=\delta_{\Gamma}^{\mathrm{BO}}+\left\|\rho\right\|$.
Definition 4.4 (Scattering matrices, [BO00, Def. 5.6]).
Let $\mathrm{Re}(\lambda)<\left\|\rho\right\|-\delta_{\Gamma}$. Then $S_{\lambda}^{\Gamma}:=\mathrm{res}_{\lambda}\circ J_{\lambda}\circ\mathrm{ext}_{\lambda}$ is called the $\lambda$-scattering matrix for $\Gamma\backslash X$.
Note that our $S_{\lambda}^{\Gamma}$ corresponds to $S_{-\lambda}$ in [BO00, Def. 5.6].
From [BO00, Lemma 5.7] we know that for $\mathrm{Re}(\lambda)<\left\|\rho\right\|-\delta_{\Gamma}$ the family
$\lambda\mapsto S_{\lambda}^{\Gamma}:C^{\pm\infty}(\Gamma\backslash\Omega,V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))\to C^{\pm\infty}(\Gamma\backslash\Omega,V_{\Gamma\backslash\Omega}(\sigma_{-\lambda}))$
is meromorphic. More precisely, we have the following result.
Proposition 4.5 (Scattering matrices, [BO00, Thm. 5.10]).
Suppose that $G\not=F_{4}^{-20}$. Then the families
$\lambda\mapsto S_{\lambda}^{\Gamma}:C^{\pm\infty}(\Gamma\backslash\Omega,V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))\to C^{\pm\infty}(\Gamma\backslash\Omega,V_{\Gamma\backslash\Omega}(\sigma_{-\lambda}))$
and
$\lambda\mapsto\mathrm{ext}_{\lambda}:C^{-\infty}(\Gamma\backslash\Omega,V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))\to{}^{\Gamma}C^{-\infty}(\partial X,V(\sigma_{\lambda})),$
originally defined for $\mathrm{Re}(\lambda)<\left\|\rho\right\|-\delta_{\Gamma}$, have meromorphic continuations to all of $\mathbb{C}$. The family $\mathrm{ext}$ has at most finite-dimensional singularities. Moreover, the following identities hold:
$\mathrm{ext}_{\lambda}$ $=$ $J_{-\lambda}\circ\mathrm{ext}_{\lambda}\circ S^{\Gamma}_{\lambda},$ (4.11)
$S^{\Gamma}_{-\lambda}\circ S^{\Gamma}_{\lambda}$ $=$ $\mathrm{id}.$ (4.12)
It is expected that the necessity of hypothesis $G\not=F_{4}^{-20}$ in Proposition 4.5 is an artefact of its proof. Note that the use of this result is the only reason why we had to restrict our attention to classical groups. We will address this issue elswhere.
Definition 4.6 (Scattering matrices for the globally symmetric space, c.f. [HHP19, §7]).
For $\zeta\in\mathcal{H}_{\mathrm{phys}}$ we define
$S_{\zeta}:=J_{i\zeta},$ (4.13)
the $\zeta$-scattering matrix for $X$.
Note that our $S_{\zeta}$ corresponds to $S_{-\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}$ in [HHP19, §7]. This takes into account that our principal series representation parameter $\lambda$ corresponds to $-\lambda$ in [HHP19], while our resolvent $R_{\zeta}$ corresponds to $R_{-\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}=R_{\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}$ in [HHP19].
5. Relating scattering matrices and resolvent kernels
The key to the comparison between scattering matrices and resolvent kernels for global symmetric spaces given in [HHP19] was the possibility to take boundary values of Laplace eigenfunctions on $X$. The boundary value maps are given as the inverses of Poisson transforms. Moreover, the proof of [BO12, Lemma 11] suggests a way how to use the knowledge about the relation between resolvent poles and scattering poles of rank one global symmetric spaces in order to determine the corresponding relation for rank one locally symmetric spaces. We will follow this strategy and start by reviewing a few facts about Poisson transforms and boundary value maps.
5.1. Poisson transforms and boundary values
We define the Poisson transform with parameter $\lambda\in{\bf\mathfrak{a}}_{\mathbb{C}}^{\ast}$ by
$\mathcal{P}_{\lambda}:C^{-\infty}(G/P;V(\sigma_{\lambda}))$ $\to C^{\infty}(G/K),$
$\mathcal{P}_{\lambda}(u)(g)$ $:=(\pi^{\lambda,-\infty}(g^{-1})u)(1_{-\lambda}),$
where $1_{-\lambda}\in C^{\infty}(G/P;V(\sigma_{-\lambda}))\cong H^{-\lambda,\infty}$ corresponds to the constant function on $K$ with value $1$ via the isomorphism (4.4). If $f\in H^{\lambda,\infty}\cong C^{\infty}(G/P;V(\sigma_{\lambda}))$ is identified with an element of $C^{-\infty}(G/P;V(\sigma_{\lambda}))$ via the injection (4.5), then
$\mathcal{P}_{\lambda}(f)(g)=\int_{K}f(gk)\mathrm{d}k=\int_{K}e^{(\lambda-\rho)(H(g^{-1}k))}f(k)\mathrm{d}k.$ (5.1)
Remark 5.1 (Comparing conventions).
Note that due to our sign convention, see Remark 4.2, our Poisson transform $\mathcal{P}_{\lambda}$ corresponds to $P_{-\lambda}^{T}$ in [BO12, § 4.2] and [BO00, Def. 4.8], respectively to $P_{-\lambda}$ in [HHP19, (2)].
For generic parameters $\lambda$ the Poisson transformations are essentially inverted by boundary value maps. In [HHP19, (29)] this reads
$\beta_{\rho-\lambda}P_{\lambda}=c(\lambda)\mathrm{Id}.$
We need to relate the boundary value maps used in [BO12, § 6.3] to the Poisson transforms used in [BO12, § 4.2] in order to relate them to the boundary value maps $\beta_{\lambda}$ used in [HHP19, Lemma 6.1]. In [BO12, § 6.3] the authors are not very precise about the construction of boundary values they are using. They simply refer to [KO77] and [OS80], but they give the set of leading exponents. We assume they are using the conventions of [Olb94], which are compatible with [BO00, BO12, HHP19]. Note that [Olb94, Thm. 4.17] (see also [Olb94, Thm. 3.9] and [Kna86, Lemma 9.28]) relates the corresponding boundary value maps with the Poisson transform in the following formula
$\beta_{1,\lambda}\circ P^{\gamma}_{\lambda}=c_{\gamma}(\lambda)\,\mathrm{Id}.$
The definitions of the $c$-functions in [HHP19] and [Olb94] are compatible. In our situation $\sigma$ and $\gamma$ are trivial, which means that the boundary value map $\beta_{1,\lambda}$ from [Olb94] equals the boundary value map $\beta_{\rho-\lambda}$ from [HHP19].
In [BO12, § 6.3] Bunke and Olbrich invoke [KO77] to guarantee the existence of boundary values for eigenfunctions of the Laplacians on the two factors of $(X\times X)\setminus D_{\Gamma}$, see [BO12, (48)]. It is defined for parameters determined by the leading exponents of Casimir eigenfunctions (see [BO12, Lemmas 10 & 11] and recall that in our situation $\sigma$ and $\tau$ are trivial).
In principle the construction of the boundary value maps is built on the study of asymptotic behavior of eigenfunctions. The construction in [KO77] depends on a specific and very technical theory, the algebraic analysis initiated by Sato. If the boundary value map were applied to moderately growing eigenfunctions on all of $X\times X$ we could use [vdBS87] to describe the boundary value maps. This description shows that the direct product situation leads to a decoupling of the asymptotics which leads to a decoupling of the boundary value map. Taking into account the notation and the various conventions from Remarks 5.1 and 4.3, the boundary value map from [BO12, § 6.3] suitably restricted is given by
$\beta_{\nu}=\beta_{\nu(\zeta)}:=\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}\otimes\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}},\qquad\zeta\in\mathcal{H}_{\mathrm{phys}},$ (5.2)
with the boundary value map of [HHP19] on the right-hand side. and $\nu\in(\mathfrak{a}_{\mathbb{C}}^{*})^{2}=\mathbb{C}^{2}$ denoting the leading exponents of the eigensections one takes boundary values of.
We restrict our attention to the non-vanishing boundary value maps. For these the parameter $\nu\in\mathbb{C}^{2}$ is given by
$\nu:=\nu(z):=(-\left\|\rho\right\|-\lambda(z),-\left\|\rho\right\|-\lambda(z)),$
where $\lambda(z):=\sqrt{\left\|\rho\right\|^{2}-z}\in\{\mathrm{Re}\,>0\}\subset\mathbb{C}$ (c.f. Remark 4.3), $z\in\mathbb{C}$ is a Casimir eigenvalue, and we use the convention from [BO12, § 6.1] that the complex vector space ${\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}$ is identified with $\mathbb{C}$ by complex linear extension of the unique order-preserving isometric isomorphism ${\bf\mathfrak{a}}^{\ast}\cong\mathbb{R}$. This means that $\rho\in{\bf\mathfrak{a}}^{\ast}$ is identified with the number $\left\|\rho\right\|\in\mathbb{R}$ (which accordingly is written simply as $\rho$ in [BO12, § 6]), where the norm is with respect to the inner product on ${\bf\mathfrak{a}}^{\ast}\cong{\bf\mathfrak{a}}$ given by restriction of the chosen bilinear form $b$ on ${\bf\mathfrak{g}}$ defining the Riemannian metric on $X=G/K$. In particular, the number $\lambda(z)\in\mathbb{C}$ corresponds to the element
$\lambda(z)=\sqrt{\left\|\rho\right\|^{2}-z}\,\frac{\rho}{\left\|\rho\right\|}=\sqrt{1-\frac{z}{\left\|\rho\right\|^{2}}}\,\rho\in{\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}.$ (5.3)
The number $-\left\|\rho\right\|-\lambda(z)$ is a leading exponent of a Casimir eigenfunction for the eigenvalue $z$. The boundary value map maps tensor products of Casimir eigenfunctions for the eigenvalue $z$ to tensor products of elements of $H^{1,\lambda(z)}$ (see [BO12, Lemmas 10 & 11] for all this and take the change of signs described in Remark 4.2 into account). Thus $\beta_{\nu}$ coincides with $\beta_{1,-\lambda(z)}\otimes\beta_{1,-\lambda(z)}$, where $\beta_{1,\lambda}$ is as in [Olb94], so that in view of (ii) we have
$\beta_{\nu}=\beta_{\rho+\lambda(z)}\otimes\beta_{\rho+\lambda(z)}$ (5.4)
with the boundary value maps from [HHP19] on the right hand side and the parameters considered as elements of ${\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}$.
To express the relation (5.4) using parameters in $\mathbb{C}$, one needs to take into account that in [HHP19] the identification of ${\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}$ with $\mathbb{C}$ differs from that in [BO12, § 6]: In [HHP19] the element $\rho\in{\bf\mathfrak{a}}^{\ast}$ gets identified with the number $\frac{\left\|\rho\right\|}{\left\|\alpha_{0}\right\|}\in\mathbb{R}$, where $\alpha_{0}\in\Sigma^{+}\subset{\bf\mathfrak{a}}^{\ast}$ is the unique reduced positive root and the norm is with respect to the Killing form. Let us choose the bilinear form $b$ in [BO12, § 6.1] such that on ${\bf\mathfrak{a}}\times{\bf\mathfrak{a}}$ it agrees with the Killing form, so that the norms on ${\bf\mathfrak{a}}_{\mathbb{C}}^{\ast}$ considered in [BO12, § 6.1] and [HHP19] agree. Furthermore, let us denote for any element $\lambda\in{\bf\mathfrak{a}}_{\mathbb{C}}^{\ast}$ the corresponding complex numbers in [BO12, § 6] and [HHP19] by $\lambda_{\mathrm{BO}}\in\mathbb{C}$ and $\lambda_{\mathrm{HHP}}\in\mathbb{C}$, respectively. Then
$\lambda_{\mathrm{BO}}=\left\|\alpha_{0}\right\|\lambda_{\mathrm{HHP}}.$ (5.5)
In particular, for the element $\lambda(z)\in{\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}$ from (5.3), which satisfies $\lambda(z)_{\mathrm{BO}}=\sqrt{\left\|\rho\right\|^{2}-z}=\sqrt{\rho_{\mathrm{BO}}^{2}-z}$, we find
$\begin{split}\lambda(z)_{\mathrm{HHP}}&=\frac{1}{\left\|\alpha_{0}\right\|}\sqrt{\left\|\rho\right\|^{2}-z}=\sqrt{\bigg{(}\frac{\left\|\rho\right\|}{\left\|\alpha_{0}\right\|}\bigg{)}^{2}-\frac{z}{\left\|\alpha_{0}\right\|^{2}}}=\sqrt{\rho_{\mathrm{HHP}}^{2}-\frac{z}{\left\|\alpha_{0}\right\|^{2}}}.\end{split}$ (5.6)
Lemma 5.2.
We have the identity of meromorphic functions
$s_{\zeta}|_{\partial X\times\partial X\setminus\mathrm{diag}_{\partial X}}=-2i\left\|\alpha_{0}\right\|\zeta\,\beta_{\nu(\zeta)}(r_{\zeta}|_{X\times X\setminus\mathrm{diag}_{X}}).$ (5.7)
Proof.
Let $f,g\in{C^{\infty}}(\partial X)$ have disjoint supports. Then applying [HHP19, last eq. on p. 20] with the parameter ${-\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}$ gives
$\left\langle s_{\zeta},f\otimes g\right\rangle$ $=\left\langle S_{\zeta}f,g\right\rangle$
$=-2i\left\|\alpha_{0}\right\|\zeta\big{<}\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}(\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}R_{\zeta})^{\prime}f,g\big{>}$
$=-2i\left\|\alpha_{0}\right\|\zeta\big{<}R_{\zeta}\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}^{\prime}f,\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}^{\prime}g\big{>}$
$=-2i\left\|\alpha_{0}\right\|\zeta\big{<}r_{\zeta}|_{X\times X\setminus\mathrm{diag}_{X}},(\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}^{\prime}\otimes\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}^{\prime})(f\otimes g)\big{>}$
$=-2i\left\|\alpha_{0}\right\|\zeta\big{<}r_{\zeta}|_{X\times X\setminus\mathrm{diag}_{X}},\beta_{\nu(\zeta)}^{\prime}(f\otimes g)\big{>}$
$=-2i\left\|\alpha_{0}\right\|\zeta\big{<}\beta_{\nu(\zeta)}(r_{\zeta}|_{X\times X\setminus\mathrm{diag}_{X}}),f\otimes g\big{>}.$
The claim follows by density. ∎
5.2. Local-global comparisons
Recall the Schwartz kernels $r_{\zeta}$ and $\tilde{r}^{\Gamma}_{\zeta}$ of $R_{\zeta}$ and $\tilde{R}^{\Gamma}_{\zeta}$, respectively from Section 3. Further, let $s_{\zeta}$ and $\tilde{s}^{\Gamma}_{\zeta}$ be the Schwartz kernels of $S_{\zeta}$ and the $\Gamma$-invariant lifts of the Schwartz kernel of $S^{\Gamma}_{i\zeta\left\|\alpha_{0}\right\|}$, respectively. We consider all of these as meromorphically extended to $\mathbb{C}$.
The proof of [BO12, Lemma 11] establishes the existence of a scalar-valued meromorphic function $\mathbb{C}\owns\zeta\mapsto a_{\zeta}$ satisfying the following formulas.
$\beta_{\nu(\zeta)}(r_{\zeta}|_{X\times X\setminus\mathrm{diag}_{X}})$ $=$ $a_{\zeta}s_{\zeta}|_{\partial X\times\partial X\setminus\mathrm{diag}_{\partial X}}$ (5.8)
$\beta_{\nu(\zeta)}((\tilde{r}^{\Gamma}_{\zeta}-r_{\zeta})|_{X\times X\setminus(D_{\Gamma}\setminus\mathrm{diag}_{X})})$ $=$ $a_{\zeta}(\tilde{s}^{\Gamma}_{\zeta}-s_{\zeta})|_{\Omega\times\Omega\setminus(D^{\Omega}_{\Gamma}\setminus\mathrm{diag}_{\Omega})},$ (5.9)
where
$D^{\Omega}_{\Gamma}:=(\{e\}\times\Gamma)\cdot\mathrm{diag}_{\Omega}=(\Gamma\times\{e\})\cdot\mathrm{diag}_{\Omega}\subset\Omega\times\Omega.$
In other words, we also have
$\beta_{\nu(\zeta)}(\tilde{r}^{\Gamma}_{\zeta}|_{X\times X\setminus D_{\Gamma}})=a_{\zeta}\tilde{s}_{\zeta}^{\Gamma}|_{\Omega\times\Omega\setminus D^{\Omega}_{\Gamma}}.$
Actually, [BO12, p. 140] does only yield Equation (5.9) on a neighborhood of $\Omega\times\Omega\setminus(D_{\Gamma}^{\Omega}\setminus\mathrm{diag}_{\Omega})$ in the closure of $X\times X$ (intersected with $X\times X\setminus\mathrm{diag}_{X})$. But this suffices as the boundary values only depend on the values on a neighborhood of the boundary.
Lemma 5.2 shows that we know the function $a_{\zeta}$ explicitly on $(\partial X\times\partial X)\setminus\mathrm{diag}_{\partial X}$:
$a_{\zeta}=\frac{i}{2\left\|\alpha_{0}\right\|\zeta}.$ (5.10)
We thus arrive at the result
$-2i\left\|\alpha_{0}\right\|\zeta\,\beta_{\nu(\zeta)}(\tilde{r}^{\Gamma}_{\zeta}|_{X\times X\setminus D_{\Gamma}})=\tilde{s}_{\zeta}^{\Gamma}|_{\Omega\times\Omega\setminus D^{\Omega}_{\Gamma}}.$ (5.11)
The following lemma will help us to remove $D_{\Gamma}$ and $D_{\Gamma}^{\Omega}$ from Equation (5.11).
Lemma 5.3.
Let $M$ be a smooth Riemannian manifold and
$P:{C^{\infty}}(M)\to{C^{\infty}}(M)$
an elliptic differential operator of order $>0$ (i.e., non-constant). Suppose that the resolvent
$(P-z)^{-1}:L^{2}(M)\to L^{2}(M)$
is defined for all $z$ in an open subset $U\subset\mathbb{C}$ and extends to a connected open subset $V\supset U$ of $\mathbb{C}$ as a meromorphic family of operators
$R(z):{C^{\infty}_{c}}(M)\to\mathcal{D}^{\prime}(M).$
Let $r(z)\in\mathcal{D}^{\prime}(M\times M)$ be the meromorphic family of distributions given by the Schwartz kernel of $R(z)$. Then a complex number $z_{0}\in V$ is a pole of $r(z)$ if, and only if, it is a pole of the meromorphic family
$r(z)|_{M\times M\setminus\mathrm{diag}_{M}}\in\mathcal{D}^{\prime}(M\times M\setminus\mathrm{diag}_{M})$
given by the restriction of $r(z)$ to the complement of the diagonal in $M\times M$.
Proof.
Denote by $P_{{C^{\infty}_{c}}}:{C^{\infty}_{c}}(M)\to{C^{\infty}_{c}}(M)$ and $P_{\mathcal{D}^{\prime}}:\mathcal{D}^{\prime}(M)\to\mathcal{D}^{\prime}(M)$ the operators obtained from $P$ by restriction and duality, respectively. Then for all $z$ in the set $U$, where $R(z)$ is $(P-z)^{-1}$ equipped with a smaller domain and a larger codomain, we have
$(P_{\mathcal{D}^{\prime}}-z\,\mathrm{Id}_{\mathcal{D}^{\prime}(M)})\circ R(z)$ $=I,$ (5.12)
$R(z)\circ(P_{{C^{\infty}_{c}}}-z\,\mathrm{Id}_{{C^{\infty}_{c}}(M)})$ $=I,$ (5.13)
where $I:{C^{\infty}_{c}}(M)\to\mathcal{D}^{\prime}(M)$ is the inclusion. Now (5.12) and (5.13) are identities of holomorphic operator families on $U$ which continue meromorphically to the connected open set $V\supset U$, so (5.12) and (5.13) also hold as identities of meromorphic operator families on $V$.
Let $z_{0}\in V$. If $z_{0}$ is a pole of $r(z)|_{M\times M\setminus\mathrm{diag}_{M}}$, then it clearly is also a pole of $r(z)$. Conversely, assume that $z_{0}$ is a pole of $r(z)$. This means that for all $z\neq z_{0}$ in a neighborhood $U_{0}$ of $z_{0}$ in $V$ we have a unique presentation of $r(z)$ as a Laurent series
$r(z)=\sum_{k\geq-k_{0}}(z-z_{0})^{k}a_{k}$ (5.14)
with distributions $a_{k}\in\mathcal{D}^{\prime}(M\times M)$ and $k_{0}\in\mathbb{N}=\{1,2,\ldots\}$ with $a_{-k_{0}}\neq 0$. Now suppose, in order to provoke a contradiction, that $z_{0}$ is not a pole of $r(z)|_{M\times M\setminus\mathrm{diag}_{M}}$. This means that for all $z$ in a neighborhood $U_{1}$ of $z_{0}$ in $V$ we have a unique presentation
$r(z)|_{M\times M\setminus\mathrm{diag}_{M}}=\sum_{k\geq 0}(z-z_{0})^{k}b_{k}$ (5.15)
with distributions $b_{k}\in\mathcal{D}^{\prime}(M\times M\setminus\mathrm{diag}_{M})$. By restricting $r(z)$ to $M\times M\setminus\mathrm{diag}_{M}$ for $z\in U_{0}\cap U_{1}$, the uniqueness of the presentation (5.15) implies that
$a_{k}|_{M\times M\setminus\mathrm{diag}_{M}}=0\qquad\forall\;k<0.$ (5.16)
For each $k\geq-k_{0}$, let $A_{k}:={C^{\infty}_{c}}(M)\to\mathcal{D}^{\prime}(M)$ be the operator with Schwartz kernel $a_{k}$. Then (5.14) implies that we have a convergent series
$R(z)=\sum_{k\geq-k_{0}}(z-z_{0})^{k}A_{k}$ (5.17)
for all $z\in U_{0}$, and (5.16) implies that $A_{k}$ is a local operator for all $k<0$: If $f\in{C^{\infty}_{c}}(M)$ and $\chi\in{C^{\infty}_{c}}(M\setminus\operatorname{supp}f)$, then (5.16) gives
$A_{k}(f)(\chi)=a_{k}(f\otimes\chi)=0$
since $\operatorname{supp}(f\otimes\chi)\subset M\times M\setminus\mathrm{diag}_{M}$. Thus $\operatorname{supp}A_{k}(f)\subset\operatorname{supp}f$.
By (5.17) we can extract the operator $A_{-k_{0}}$ from the meromorphic family $R(z)$ as
$A_{-k_{0}}=\lim_{\begin{subarray}{c}z\to z_{0}\\ z\neq z_{0}\end{subarray}}(z-z_{0})^{k_{0}}R(z).$
On the other hand, by (5.12), we have
$(P_{\mathcal{D}^{\prime}}-z\,\mathrm{Id}_{\mathcal{D}^{\prime}(M)})\circ(z-z_{0})^{k_{0}}R(z)$ $=(z-z_{0})^{k_{0}}(P_{\mathcal{D}^{\prime}}-z\,\mathrm{Id}_{\mathcal{D}^{\prime}(M)})\circ R(z)$
$=(z-z_{0})^{k_{0}}I$
$\stackrel{{\scriptstyle z\to z_{0}}}{{\longrightarrow}}0.$
Using the equicontinuity of the operator family $P_{\mathcal{D}^{\prime}}-z\,\mathrm{Id}_{\mathcal{D}^{\prime}(M)}$ at $z=z_{0}$ we deduce that
$(P_{\mathcal{D}^{\prime}}-z_{0}\,\mathrm{Id}_{\mathcal{D}^{\prime}(M)})\circ A_{-k_{0}}=0,$ (5.18)
i.e.,
$\mathrm{im}(A_{-k_{0}})\subset\ker(P_{\mathcal{D}^{\prime}}-z_{0}\,\mathrm{Id}_{\mathcal{D}^{\prime}(M)}).$ (5.19)
Now $P$ is elliptic, so (5.19) gives $\mathrm{im}(A_{-k_{0}})\subset{C^{\infty}}(M)$ by elliptic regularity. We already know that $A_{-k_{0}}$ is a local operator, so by Peetre’s Theorem we conclude that $A_{-k_{0}}$ is a differential operator. In particular, $\mathrm{im}(A_{-k_{0}})\subset{C^{\infty}_{c}}(M)$. Knowing this, (5.18) implies that
$(P_{{C^{\infty}_{c}}}-z_{0}\,\mathrm{Id}_{{C^{\infty}_{c}}(M)})\circ A_{-k_{0}}=0.$ (5.20)
Now, around each $x\in M$ we can choose a neighborhood $W$ of $x$ in $M$ such that the differential operator $A_{-k_{0}}$ has a constant order $N$ on $W$. Let $\sigma_{P},\sigma_{A_{-k_{0}}}\in{C^{\infty}}(T^{\ast}M)$ be the principal symbols of $P$ and $A_{-k_{0}}$, respectively. By the ellipticity of $P$ we have $\sigma_{P}(x,\xi)\neq 0$ for all $\xi\neq 0$. Since $P$ has order $>0$ it follows that $\sigma_{P}(x,\xi)-z_{0}\neq 0$ for all $\xi$ with $\left\|\xi\right\|$ sufficiently large. From the composition theorem and (5.20) we get
$(\sigma_{P}(x,\xi)-z_{0})\sigma_{A_{-k_{0}}}=0,$
which implies that $\sigma_{A_{-k_{0}}}(x,\xi)=0$ for all $\xi$ with $\left\|\xi\right\|$ sufficiently large and consequently $\sigma_{A_{-k_{0}}}(x,\xi)=0$ for all $\xi$ because it is a polynomial. Thus the order $N$ component of $A_{-k_{0}}$ vanishes at $x$, which implies that $A_{-k_{0}}|_{W}=0$. Since $x\in M$ was arbitrary, we conclude that $A_{-k_{0}}=0$ and hence $a_{-k_{0}}=0$, a contradiction. ∎
We can now prove our main result:
Theorem 5.4.
Outside the set $-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}$, the scattering poles which are not in the physical plane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ are precisely the quantum resonances.
Proof.
Let $\zeta_{0}$ be a quantum resonance of $\Gamma\backslash X$, i.e. a pole of $\tilde{r}_{\zeta}^{\Gamma}$. Then $\zeta_{0}$ is not in the physical halfplane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ and Lemma 5.3 implies that $\zeta_{0}$ is also a pole of $\tilde{r}_{\zeta}^{\Gamma}|_{X\times X\backslash D_{\Gamma}}$.
Case 1 $\zeta_{0}\not\in-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}$ and $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}\not=0$:
In this case Equation (5.2) and the result [HHP19, Thm. 6.4] imply that $\zeta\mapsto\beta_{\nu(\zeta)}$ is a holomorphic family of injective operators in a neighborhood of $\zeta_{0}$. Thus by Equation (5.11) $\zeta_{0}$ is a pole of $\tilde{s}_{\zeta}^{\Gamma}|_{\Omega\times\Omega\setminus D^{\Omega}_{\Gamma}}$. Hence $\zeta_{0}$ is also a pole of $\tilde{s}_{\zeta}^{\Gamma}$, i.e. a scattering pole of $\Gamma\backslash X$.
Case 2 $\zeta_{0}\not\in-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}$ and $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}=0$:
In view of [HHP19, Re. 5.1] $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}=0$ implies $\mathrm{Im}\,\zeta_{0}>\|\rho\|$. If $\delta_{\Gamma}<2\|\rho\|$, then $i\rho\in\mathcal{H}^{\Gamma}_{\mathrm{phys}}$, hence also $\zeta_{0}\in\mathcal{H}^{\Gamma}_{\mathrm{phys}}$. Thus, this case cannot occur, i.e. $\delta_{\Gamma}=2\|\rho\|$. But then 00 is in the $L^{2}$-spectrum of $\Delta_{\Gamma\backslash X}$. Since the continuous $L^{2}$-spectrum of $\Delta_{\Gamma\backslash X}$ is $[\|\rho\|^{2},\infty[$ the point 00 has to be an eigenvalue. But $\Delta_{\Gamma\backslash X}f=0$ for $f\in L^{2}(\Gamma\backslash X)$ is equivalent to $df=0$ and $d^{*}f=0$, so $f$ has to be constant. But this implies that $\Gamma\backslash X$ has finite volume, which contradicts our hypothesis the $\Gamma$ is convex-cocompact, but not cocompact.
Conversely, assume that $\zeta_{0}$ is a scattering pole which is not in the physical halfplane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ and satisfies $\zeta_{0}\not\in-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}$. Then, by the argument given in Case 2, $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}\not=0$ and again Equation (5.2) and [HHP19, Thm. 6.4] imply that $\zeta\mapsto\beta_{\nu(\zeta)}$ is a holomorphic family of injective operators in a neighborhood of $\zeta_{0}$.
For any $\lambda\in{\bf\mathfrak{a}}_{\mathbb{C}}^{\ast}$ and $f\in C^{-\infty}(\Gamma\backslash\Omega,V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ [BO12, p. 134] yields an Eisenstein series, which we denote by $\mathcal{P}_{\lambda}^{\Gamma}f$ since it takes the role of the Poisson transform in the case of locally symmetric spaces. Then $\mathcal{P}_{\lambda}^{\Gamma}f\in C^{\infty}(\Gamma\backslash X)$ is a $\Delta_{\Gamma\backslash X}$-eigenfunction, which has $f$ as boundary value. Applying this to the product $\Gamma\backslash\Omega\times\Gamma\backslash\Omega$ and scattering kernel $\tilde{s}_{\zeta}^{\Gamma}$ with the appropriate $\lambda$ determined by $\zeta$ we find a kernel function $r^{\Gamma,\mathrm{Eis}}_{\zeta}$ on $\Gamma\backslash X\times\Gamma\backslash X$ which has the same boundary values as $r^{\Gamma}_{\zeta}$. But then, since $\zeta\mapsto\beta_{\nu(\zeta)}$ in a neighborhood of $\zeta_{0}$ is a holomorphic family of injective maps, we have $r^{\Gamma,\mathrm{Eis}}_{\zeta}=r^{\Gamma}_{\zeta}$ and $\zeta_{0}$ is a pole of $r^{\Gamma}_{\zeta}$, i.e. it is also a quantum resonance. ∎